# Cardiocirculatory Computational Models for the Study of Hypertension

### v1

@@ 1 Introduction
Hypertension, along with higher than normal or elevated blood pressure levels, is one of the world’s leading risk factors for death and disability [1, 2], since it may cause stroke, ischaemic heart disease, altered vascular conditions and kidney disorders [3, 4]. High systolic blood pressure is responsible for more deaths than any other health risk. According to [1], it accounted for over 10 million deaths annually and contributed to over half of all cardiocirculatory-related deaths worldwide in 2019 [5]; furthermore, in less than 30 years, the number of hypertensive individuals doubled: from 650 million in 1990 to 1.3 billion in 2019 [6]. According to the most recent World Health Organization report [2] “in 2024, an estimated 1.4 billion people aged 30–79 years were affected worldwide, yet fewer than 1 in 5 (320 million) had the condition adequately controlled”.
@@
Hypertension is often referred to as the “silent killer” because it is typically asymptomatic: unless blood pressure is measured, most individuals remain unaware of their condition until they experience a clinical complication, i.e. heart attack, stroke or kidney failure. Proper management of hypertension requires accurate diagnosis, prompt initiation of treatment in severe cases and continuous monitoring of progress. Timelines are crucial: the longer a person suffers from undiagnosed or poorly managed hypertension, the worse the health consequences are likely to be [1, 2].
@@
Due to hypertension profound impact on global health, a comprehensive understanding of this condition is crucial. Mathematical models are increasingly recognized as indispensable tools for studying complex physiological phenomena, offering insights that traditional approaches alone may not provide. Within this context, the concept of Digital Twins (DTs) is gaining ground in cardiovascular research [7, 8]. Unlike a simple patient-specific model, a digital twin establishes a continuous interaction between the physical and digital entities: patient data drive the calibration of the computational model, while the predictions obtained from the model can guide clinical choices, such as the selection of therapeutic strategies [9]. By integrating clinical and sensor data into mathematical models, DTs allow simulations tailored to the individual patient, enabling accurate reproduction of haemodynamics and prediction of disease progression. Hypertension has already been investigated using machine learning models, as evidenced by works such as [10, 11]; nonetheless, there has also been an increasing use of differential equation (DE)-based models to simulate the cardiocirculatory system and study hypertension (e.g. [12, 13]). The use of DE-based models, combined with real patient data, enables the creation of personalized models tailored to individual patients: by incorporating patient-specific physiological information, these models are able to simulate cardiocirculatory haemodynamics with greater accuracy and reflect the unique characteristics of each patient’s condition. When combined with real patient data, such DE-based models naturally lend themselves to the construction of DTs, as they can reproduce patient-specific cardiocirculatory dynamics and reflect the unique characteristics of individual conditions.
@@
Several differential mathematical and numerical models are employed to analyze the dynamics of hypertension, simulate its effects under various conditions and compare the simulated outputs with clinically observed trends and reference values from the literature: in particular, a lumped-parameter cardiocirculatory model (from now on referred to as $(\mathscr{C})$ or 0D model), based on a 0D representation by an RLC circuit, and an electromechanical model ([14]) coupled with a lumped-parameter model (from now on referred to as 3D–0D model), where the left ventricle is represented in 3D through sophisticated electromechanical modeling, coupled with a 0D model of the blood circulation. The integrated approach not only enhances the understanding of hypertension, but also contributes to the development of improved diagnostic, therapeutic and preventive strategies, within a DT framework. Through the proposed modeling framework, the simulations reproduce the haemodynamic alterations associated with various forms and grades of hypertension; the results show the effect of increased vascular resistance and reduced compliance on both systemic and pulmonary circulation, demonstrating how these changes deeply affect pressure dynamics and cardiac workload. Furthermore, the comparison between 0D and 3D–0D models yields valuable insights into advantages and limitations of each approach, emphasizing the importance of patient-specific modeling for an accurate representation of hypertensive conditions.
@@ 2 Pathophysiology of Hypertension
In accordance with the 2023 European and current international guidelines [15], hypertension is diagnosed when office blood pressure measurements (i.e. taken in a clinical or healthcare setting) is equal or higher than 140 mmHg for systolic blood pressure or 90 mmHg for diastolic blood pressure, although recent USA guidelines have adopted a lower threshold, i.e. equal or higher than 130/80 mmHg. Alternative definitions are considered to identify related conditions such as pulmonary hypertension, as described below. Hypertension can be classified in several ways considering aspects such as the underlying causes, the severity of blood pressure elevation, the timing and consistency of measurements, the presence of associated complications, or other risk factors. An important distinction is made between primary and secondary hypertension: while primary hypertension has no clearly identifiable cause and accounts for the vast majority of cases, secondary hypertension arises from specific, diagnosable medical conditions (e.g. renal, endocrine or vascular disorders) [16]. Hypertension can be further classified based on its origin and the vascular district involved. Here, the forms that are more related to this work are briefly described:
@@
    systemic hypertension refers to elevated pressure in the systemic circulation and is the most common form [16]. It is typically associated with increased peripheral resistance and large artery stiffening, affecting left ventricular function and contributing to long-term cardiovascular damage;
@@
    pulmonary hypertension by contrast, is defined as a mean pulmonary artery pressure above 20 mmHg at rest [17]. It affects the pulmonary circulation and can be caused by a variety of mechanisms, including left heart dysfunction, lung diseases, hypoxia exposure, or vascular abnormalities. It leads to increased right ventricular workload and is associated with a high morbidity and mortality risk;
@@
    renovascular hypertension is a form of secondary hypertension caused by reduced renal perfusion, most commonly due to atherosclerotic renal artery stenosis or fibromuscular dysplasia [18]. The decrease in renal blood flow promotes renin secretion, vasoconstriction, sympathetic stimulation and fluid retention, all of which contribute to a sustained elevation in systemic blood pressure.
@@ 2.1 Etiology and Clinical Manifestations
Hypertension is a multifactorial condition with both genetic and environmental influences. It may be asymptomatic in its early stages but can lead to progressive damage to key organs such as the heart, brain, vasculature and kidneys as blood pressure rises [19]. Hypertensive individuals often exhibit enhanced vasoconstriction in both large and small arteries. This phenomenon, alongside the increase in blood pressure, appears to be favored by systemic autoregulation, triggered by the preliminary increase of blood volume and cardiac output; this mechanism refers to the ability of blood vessels to maintain a constant blood flow, despite changes in blood pressure [20]. Hypertension is very commonly associated with impaired vasodilation, due mostly to endothelial dysfunction and structural changes in blood vessels [21]. In small vessels, remodeling of arterioles results in smooth muscle layer thickening and lumen narrowing, increasing overall vascular resistance [22]. Other studies documented a two-way relation between aortic stiffness and hypertension, as one condition increases due to the other one in a self-promoting circle. A known impact of stiffness is the transmission of pressure into the smaller vessels, which may cause harm to the organs [23].
Chronic high blood pressure puts stress on the left ventricular wall by increasing the afterload, promoting left ventricular remodeling and hypertrophy; the latter is a well-established risk factor for adverse cardiocirculatory outcomes, such as myocardial infarction, heart failure, atrial fibrillation, stroke and cardiocirculatory mortality [19]. Long-term elevated pressure in the left ventricle may lead to the dysfunction of the left atrium through dilation. There is ample evidence linking hypertension with the etiology of valvular heart disease [24]: an increased systolic blood pressure has been associated with an elevated risk of conditions such as aortic stenosis and aortic regurgitation. The effects of hypertension involve endothelial dysfunction, vascular remodeling and enhanced vascular stiffness, which not only contribute to the development of the disease, but are also results of it in turn [25]. The kidney is also a key actor in the development and sustainment of hypertension: regulating systemic vasoconstriction and blood pressure through both direct hormonal actions and feedback mechanisms, including fluid balance, sodium handling and sympathetic activity [26, 27]. Hypertension and chronic kidney disease are closely linked in a complex relationship: the former can be considered both a cause and a consequence of the latter, promoting its progression [28].
@@ 3 Differential Mathematical Models for the Cardiocirculatory System
This section provides an overview of the models adopted for this paper, describing their mathematical formulation, the key assumptions, the functionality and the involved parameters [29, 30, 31]. In Section 3.1, the 0D models are illustrated, while the 3D–0D coupled model is described and briefly analyzed in Section 3.2.
@@ 3.1 A 0D Model for Closed-Loop Blood Circulation
The 0D model $(\mathscr{C})$ [29, 31] describes the haemodynamics of the cardiocirculatory system by means of a lumped-parameter approach: it represents the entire circulatory network as a set of 0D interconnected components, and allows the system to be modeled using ordinary differential equations (ODEs). $(\mathscr{C})$ is based on the assumption that the cardiocirculatory system can be simplified into discrete elements that capture the mean features of blood flow and pressure dynamics. This type of models have been implemented in two variants (Figure LABEL:fig:lumped_nocap): $(\mathscr{C}_{\text{C}})$ which includes capillaries and $(\mathscr{C}_{\text{NC}})$ which does not. The circulation is simulated over a time period T, long enough to reach a periodic limit cycle in the output variables, with period that of the cardiac cycle; this periodicity is a property of the simulation rather than the model itself. Only results belonging to the last heartbeat, lasting $\text{T}_{\text{HB}}$, are then considered.
@@
Figure 1: 0D cardiocirculatory models with (center) and without (left) capillaries. The 3D–0D coupling between the left ventricle 3D electromechanical model and the 0D circulation model without capillaries is shown on the right. Source: [29, 31] and graphical modifications.
@@ 3.2 The 3D Electromechanical Model coupled with the Lumped-Parameter Model
The 3D–0D model employed in this paper is the one presented in [29], which couples a detailed 3D electromechanical model of the left ventricle with a 0D model of the systemic and pulmonary circulations. The 3D left ventricle is represented by the computational domain $\Omega_{0}\subset\mathbb{R}^{3}$, with its boundary $\partial\Omega_{0}$ being split into three distinct sections: the endocardium ($\Gamma^{\text{endo}}_{0}$), the epicardium ($\Gamma^{\text{epi}}_{0}$) and the ventricular base ($\Gamma^{\text{base}}_{0}$); the latter represents the artificial boundary where the left ventricle geometry is cut.
@@
The 3D–0D complete model is defined by sets of partial differential equations (PDEs); the detailed formulation is reported in A. In $(\mathscr{E})$, the dynamics of the transmembrane potential $u$ are captured by a diffusion-reaction PDE, which governs the electrical behaviour of cardiac muscle cells [32, 33, 34]. For the purpose of electromechanic coupling, a proper modeling of the ionic fluxes across the cell membrane is mandatory [35]; the latter involves the inclusion of two variables: the recovery variables $\boldsymbol{w}$, which characterize the proportion of open ionic channels, and the concentration variables $\boldsymbol{z}$, describing the concentrations of critical ionic species, such as intracellular calcium ions $\left[\text{Ca}^{2+}\right]_{i}$, which play a pivotal role in triggering mechanical contraction.
@@
In the electrophysiological framework defined by $(\mathscr{E})-(\mathscr{I})$, the transmembrane potential $u$ propagates through gap junctions; this phenomenon is governed by the diffusion term $\nabla\cdot\left(J\mathbf{F}^{-1}\mathbf{D}_{\text{M}}\mathbf{F}^{-T}\nabla u\right)$, which incorporates the influence of tissue deformation and introduces a feedback mechanism, linking electrical behaviour to mechanical stretch. Here, $\mathbf{F}=\mathbf{I}+\nabla\boldsymbol{d}$ is the deformation gradient tensor, while $\mathbf{D}_{\text{M}}$ represents the diffusion tensor, which captures the anisotropic characteristics of cardiac tissue. The applied current, $\mathcal{I}_{\text{app}}(t)$, simulates the role of the Purkinje network: these fibers form the outermost component of the cardiac conduction system and their primary function is to ensure the swift and synchronized activation of the ventricular myocardium [36]. Additionally, the ionic current $\mathcal{I}_{\text{ion}}(u,\boldsymbol{w},\boldsymbol{z})$ encapsulates multiscale interactions at both cellular and tissue levels. Electrically isolated interfaces are modeled using homogeneous Neumann boundary conditions.
@@
Cardiac contraction arises from interactions between actin and myosin proteins within sarcomeres, the fundamental contractile units of heart muscle [37]. To simulate the mechanism in $(\mathscr{A})$, the RDQ20-MF model has been employed [38] instead of RDQ18 in [29]: it provides a complete biophysical model that characterizes the behaviour of sarcomeric proteins, along with the causes that support the responsiveness of the heart towards variations in calcium levels. RDQ20-MF is a system of differential equations, where the state variables are represented by the vector $\boldsymbol{s}$; it outputs the permissivity $P$, representing the fraction of contractile units in the force-generating state, from which the effective active tension $T_{a}$ is derived. Inputs to the model include calcium concentration $\left[\text{Ca}^{2+}\right]_{i}$ (properly described using a ionic model), sarcomere length (SL) and its time derivative from the mechanical model.
@@
Tissue displacement $\boldsymbol{d}$ is described by the momentum conservation equation $(\mathscr{M})$. The Piola-Kirchhoff stress tensor $\mathbf{P}$ represents both the passive and active mechanical properties of the tissue. Under the assumption of hyperelasticity, the passive component of the stress tensor is derived from the strain energy density function $\mathcal{W}$, while the active component depends on tissue stretch along the fiber direction and on the active tension $T_{a}$. To capture the anisotropic properties of cardiac muscle, the Guccione strain energy density function as $\mathcal{W}(\mathbf{F})$ [39, 40] is used, which depends on the Green-Lagrange strain energy tensor. Additionally, to penalize large volume variations and enforce weak incompressibility, a term involving $J=\det\mathbf{F}$ is added to $\mathcal{W}(\mathbf{F})$ [41].
@@
To model the interaction between the left ventricle and the pericardium [42], a generalized Robin boundary condition is applied on the epicardium $\Gamma_{0}^{\text{epi}}$: this condition is designed to simulate the effect of the right ventricle, major veins and arteries, i.e. to restrict rigid rotations of the ventricle around the apico-basal axis (the direction from the base of the heart to its apex), while permitting torsion. At the base $\Gamma^{\text{base}}_{0}$, an energy-consistent boundary condition is enforced: this ensures the correct stress distribution at the base boundary. For the endocardium $\Gamma^{\text{endo}}_{0}$, the applied boundary condition accounts for the pressure $p_{\text{LV}}(t)$, which reflects the pressure exerted by the blood within the ventricle. Lastly, the mechanical model $(\mathscr{M})$ influences $(\mathscr{A})$ by determining the local sarcomere length (SL); the sarcomeres are aligned with the muscle fibers, so the local sarcomere length directly correlates with the tissue stretch along the fiber direction.

### v2

@@ 1 Introduction
Hypertension, along with higher than normal or elevated blood pressure levels, is one of the world’s leading risk factors for death and disability [who2023, who2025], since it may cause stroke, ischaemic heart disease, altered vascular conditions and kidney disorders [intro:1, intro:2]. High systolic blood pressure is responsible for more deaths than any other health risk. According to [who2023], it accounted for over 10 million deaths annually and contributed to over half of all cardiocirculatory-related deaths worldwide in 2019 [intro:5]; furthermore, in less than 30 years, the number of hypertensive individuals doubled: from 650 million in 1990 to 1.3 billion in 2019 [intro:3]. According to the most recent World Health Organization report [who2025] “in 2024, an estimated 1.4 billion people aged 30–79 years were affected worldwide, yet fewer than 1 in 5 (320 million) had the condition adequately controlled”.
@@
Hypertension is often referred to as the “silent killer” because it is typically asymptomatic: unless blood pressure is measured, most individuals remain unaware of their condition until they experience a clinical complication, i.e. heart attack, stroke or kidney failure. Proper management of hypertension requires accurate diagnosis, prompt initiation of treatment in severe cases and continuous monitoring of progress. Timelines are crucial: the longer a person suffers from undiagnosed or poorly managed hypertension, the worse the health consequences are likely to be [who2023, who2025].
@@
Due to hypertension profound impact on global health, a comprehensive understanding of this condition is crucial. Mathematical models are increasingly recognized as indispensable tools for studying complex physiological phenomena, offering insights that traditional approaches alone may not provide. Within this context, the concept of Digital Twins (DTs) is gaining ground in cardiovascular research [dt2, dt3]. Unlike a simple patient-specific model, a digital twin establishes a continuous interaction between the physical and digital entities: patient data drive the calibration of the computational model, while the predictions obtained from the model can guide clinical choices, such as the selection of therapeutic strategies [dt1]. By integrating clinical and sensor data into mathematical models, DTs allow simulations tailored to the individual patient, enabling accurate reproduction of haemodynamics and prediction of disease progression. Hypertension has already been investigated using machine learning models, as evidenced by works such as [es:mach_learn1, es:mach_learn2]; nonetheless, there has also been an increasing use of differential equation (DE)-based models to simulate the cardiocirculatory system and study hypertension (e.g. [model_hyper:es, es:hyp_study]). The use of DE-based models, combined with real patient data, enables the creation of personalized models tailored to individual patients: by incorporating patient-specific physiological information, these models are able to simulate cardiocirculatory haemodynamics with greater accuracy and reflect the unique characteristics of each patient’s condition. When combined with real patient data, such DE-based models naturally lend themselves to the construction of DTs, as they can reproduce patient-specific cardiocirculatory dynamics and reflect the unique characteristics of individual conditions.
@@
Several differential mathematical and numerical models are employed to analyze the dynamics of hypertension, simulate its effects under various conditions and compare the simulated outputs with clinically observed trends and reference values from the literature: in particular, a lumped-parameter cardiocirculatory model (from now on referred to as $(\mathscr{C})$ or 0D model), based on a 0D representation by an RLC circuit, and an electromechanical model ([FEDELE2023115983]) coupled with a lumped-parameter model (from now on referred to as 3D–0D model), where the left ventricle is represented in 3D through sophisticated electromechanical modeling, coupled with a 0D model of the blood circulation. The integrated approach not only enhances the understanding of hypertension, but also contributes to the development of improved diagnostic, therapeutic and preventive strategies, within a DT framework. Through the proposed modeling framework, the simulations reproduce the haemodynamic alterations associated with various forms and grades of hypertension; the results show the effect of increased vascular resistance and reduced compliance on both systemic and pulmonary circulation, demonstrating how these changes deeply affect pressure dynamics and cardiac workload. Furthermore, the comparison between 0D and 3D–0D models yields valuable insights into advantages and limitations of each approach, emphasizing the importance of patient-specific modeling for an accurate representation of hypertensive conditions.
@@ 2 Pathophysiology of Hypertension
In accordance with the 2023 European and current international guidelines [guidelines:eu], hypertension is diagnosed when office blood pressure measurements (i.e. taken in a clinical or healthcare setting) is equal or higher than 140 mmHg for systolic blood pressure or 90 mmHg for diastolic blood pressure, although recent USA guidelines have adopted a lower threshold, i.e. equal or higher than 130/80 mmHg. Alternative definitions are considered to identify related conditions such as pulmonary hypertension, as described below. Hypertension can be classified in several ways considering aspects such as the underlying causes, the severity of blood pressure elevation, the timing and consistency of measurements, the presence of associated complications, or other risk factors. An important distinction is made between primary and secondary hypertension: while primary hypertension has no clearly identifiable cause and accounts for the vast majority of cases, secondary hypertension arises from specific, diagnosable medical conditions (e.g. renal, endocrine or vascular disorders) [percent_esshyper]. Hypertension can be further classified based on its origin and the vascular district involved. Here, the forms that are more related to this work are briefly described:
@@
    systemic hypertension refers to elevated pressure in the systemic circulation and is the most common form [percent_esshyper]. It is typically associated with increased peripheral resistance and large artery stiffening, affecting left ventricular function and contributing to long-term cardiovascular damage;
@@
    pulmonary hypertension by contrast, is defined as a mean pulmonary artery pressure above 20 mmHg at rest [pulhyper:guidelines]. It affects the pulmonary circulation and can be caused by a variety of mechanisms, including left heart dysfunction, lung diseases, hypoxia exposure, or vascular abnormalities. It leads to increased right ventricular workload and is associated with a high morbidity and mortality risk;
@@
    renovascular hypertension is a form of secondary hypertension caused by reduced renal perfusion, most commonly due to atherosclerotic renal artery stenosis or fibromuscular dysplasia [reno-def]. The decrease in renal blood flow promotes renin secretion, vasoconstriction, sympathetic stimulation and fluid retention, all of which contribute to a sustained elevation in systemic blood pressure.
@@ 2.1 Etiology and Clinical Manifestations
Hypertension is a multifactorial condition with both genetic and environmental influences. It may be asymptomatic in its early stages but can lead to progressive damage to key organs such as the heart, brain, vasculature and kidneys as blood pressure rises [guidelines:2024esc]. Hypertensive individuals often exhibit enhanced vasoconstriction in both large and small arteries. This phenomenon, alongside the increase in blood pressure, appears to be favored by systemic autoregulation, triggered by the preliminary increase of blood volume and cardiac output; this mechanism refers to the ability of blood vessels to maintain a constant blood flow, despite changes in blood pressure [autoregulation]. Hypertension is very commonly associated with impaired vasodilation, due mostly to endothelial dysfunction and structural changes in blood vessels [vasc5]. In small vessels, remodeling of arterioles results in smooth muscle layer thickening and lumen narrowing, increasing overall vascular resistance [vasc11]. Other studies documented a two-way relation between aortic stiffness and hypertension, as one condition increases due to the other one in a self-promoting circle. A known impact of stiffness is the transmission of pressure into the smaller vessels, which may cause harm to the organs [vasc20].
Chronic high blood pressure puts stress on the left ventricular wall by increasing the afterload, promoting left ventricular remodeling and hypertrophy; the latter is a well-established risk factor for adverse cardiocirculatory outcomes, such as myocardial infarction, heart failure, atrial fibrillation, stroke and cardiocirculatory mortality [guidelines:2024esc]. Long-term elevated pressure in the left ventricle may lead to the dysfunction of the left atrium through dilation. There is ample evidence linking hypertension with the etiology of valvular heart disease [symptoms:heart14]: an increased systolic blood pressure has been associated with an elevated risk of conditions such as aortic stenosis and aortic regurgitation. The effects of hypertension involve endothelial dysfunction, vascular remodeling and enhanced vascular stiffness, which not only contribute to the development of the disease, but are also results of it in turn [symptoms:heart17]. The kidney is also a key actor in the development and sustainment of hypertension: regulating systemic vasoconstriction and blood pressure through both direct hormonal actions and feedback mechanisms, including fluid balance, sodium handling and sympathetic activity [kidney6, kidney8]. Hypertension and chronic kidney disease are closely linked in a complex relationship: the former can be considered both a cause and a consequence of the latter, promoting its progression [symptoms:kidney2].
@@ 3 Differential Mathematical Models for the Cardiocirculatory System
This section provides an overview of the models adopted for this paper, describing their mathematical formulation, the key assumptions, the functionality and the involved parameters [3d-0d_ventr, model-whole-heart1, lumped:capillary]. In Section 3.1, the 0D models are illustrated, while the 3D–0D coupled model is described and briefly analyzed in Section 3.2.
@@ 3.1 A 0D Model for Closed-Loop Blood Circulation
The 0D model $(\mathscr{C})$ [3d-0d_ventr, lumped:capillary] describes the haemodynamics of the cardiocirculatory system by means of a lumped-parameter approach: it represents the entire circulatory network as a set of 0D interconnected components, and allows the system to be modeled using ordinary differential equations (ODEs). $(\mathscr{C})$ is based on the assumption that the cardiocirculatory system can be simplified into discrete elements that capture the mean features of blood flow and pressure dynamics. This type of models have been implemented in two variants (Figure LABEL:fig:lumped_nocap): $(\mathscr{C}_{\text{C}})$ which includes capillaries and $(\mathscr{C}_{\text{NC}})$ which does not. The circulation is simulated over a time period T, long enough to reach a periodic limit cycle in the output variables, with period that of the cardiac cycle; this periodicity is a property of the simulation rather than the model itself. Only results belonging to the last heartbeat, lasting $\text{T}_{\text{HB}}$, are then considered.
@@
Figure 1: 0D cardiocirculatory models with (center) and without (left) capillaries. The 3D–0D coupling between the left ventricle 3D electromechanical model and the 0D circulation model without capillaries is shown on the right. Source: [3d-0d_ventr, lumped:capillary] and graphical modifications.
@@ 3.2 The 3D Electromechanical Model coupled with the Lumped-Parameter Model
The 3D–0D model employed in this paper is the one presented in [3d-0d_ventr], which couples a detailed 3D electromechanical model of the left ventricle with a 0D model of the systemic and pulmonary circulations. The 3D left ventricle is represented by the computational domain $\Omega_{0}\subset\mathbb{R}^{3}$, with its boundary $\partial\Omega_{0}$ being split into three distinct sections: the endocardium ($\Gamma^{\text{endo}}_{0}$), the epicardium ($\Gamma^{\text{epi}}_{0}$) and the ventricular base ($\Gamma^{\text{base}}_{0}$); the latter represents the artificial boundary where the left ventricle geometry is cut.
@@
The 3D–0D complete model is defined by sets of partial differential equations (PDEs); the detailed formulation is reported in A. In $(\mathscr{E})$, the dynamics of the transmembrane potential $u$ are captured by a diffusion-reaction PDE, which governs the electrical behaviour of cardiac muscle cells [3d-0d:model-e:behaviour1, 3d-0d:model-e:behaviour2, 3d-0d:model-e:behaviour3]. For the purpose of electromechanic coupling, a proper modeling of the ionic fluxes across the cell membrane is mandatory [3d-0d:model-e:book]; the latter involves the inclusion of two variables: the recovery variables $\boldsymbol{w}$, which characterize the proportion of open ionic channels, and the concentration variables $\boldsymbol{z}$, describing the concentrations of critical ionic species, such as intracellular calcium ions $\left[\text{Ca}^{2+}\right]_{i}$, which play a pivotal role in triggering mechanical contraction.
@@
In the electrophysiological framework defined by $(\mathscr{E})-(\mathscr{I})$, the transmembrane potential $u$ propagates through gap junctions; this phenomenon is governed by the diffusion term $\nabla\cdot\left(J\mathbf{F}^{-1}\mathbf{D}_{\text{M}}\mathbf{F}^{-T}\nabla u\right)$, which incorporates the influence of tissue deformation and introduces a feedback mechanism, linking electrical behaviour to mechanical stretch. Here, $\mathbf{F}=\mathbf{I}+\nabla\boldsymbol{d}$ is the deformation gradient tensor, while $\mathbf{D}_{\text{M}}$ represents the diffusion tensor, which captures the anisotropic characteristics of cardiac tissue. The applied current, $\mathcal{I}_{\text{app}}(t)$, simulates the role of the Purkinje network: these fibers form the outermost component of the cardiac conduction system and their primary function is to ensure the swift and synchronized activation of the ventricular myocardium [purkinje]. Additionally, the ionic current $\mathcal{I}_{\text{ion}}(u,\boldsymbol{w},\boldsymbol{z})$ encapsulates multiscale interactions at both cellular and tissue levels. Electrically isolated interfaces are modeled using homogeneous Neumann boundary conditions.
@@
Cardiac contraction arises from interactions between actin and myosin proteins within sarcomeres, the fundamental contractile units of heart muscle [3d-0d:model-a]. To simulate the mechanism in $(\mathscr{A})$, the RDQ20-MF model has been employed [RDQ20] instead of RDQ18 in [3d-0d_ventr]: it provides a complete biophysical model that characterizes the behaviour of sarcomeric proteins, along with the causes that support the responsiveness of the heart towards variations in calcium levels. RDQ20-MF is a system of differential equations, where the state variables are represented by the vector $\boldsymbol{s}$; it outputs the permissivity $P$, representing the fraction of contractile units in the force-generating state, from which the effective active tension $T_{a}$ is derived. Inputs to the model include calcium concentration $\left[\text{Ca}^{2+}\right]_{i}$ (properly described using a ionic model), sarcomere length (SL) and its time derivative from the mechanical model.
@@
Tissue displacement $\boldsymbol{d}$ is described by the momentum conservation equation $(\mathscr{M})$. The Piola-Kirchhoff stress tensor $\mathbf{P}$ represents both the passive and active mechanical properties of the tissue. Under the assumption of hyperelasticity, the passive component of the stress tensor is derived from the strain energy density function $\mathcal{W}$, while the active component depends on tissue stretch along the fiber direction and on the active tension $T_{a}$. To capture the anisotropic properties of cardiac muscle, the Guccione strain energy density function as $\mathcal{W}(\mathbf{F})$ [guccione1, guccione2] is used, which depends on the Green-Lagrange strain energy tensor. Additionally, to penalize large volume variations and enforce weak incompressibility, a term involving $J=\det\mathbf{F}$ is added to $\mathcal{W}(\mathbf{F})$ [3d-0d:model-m].
@@
To model the interaction between the left ventricle and the pericardium [3d-0d:model-m2], a generalized Robin boundary condition is applied on the epicardium $\Gamma_{0}^{\text{epi}}$: this condition is designed to simulate the effect of the right ventricle, major veins and arteries, i.e. to restrict rigid rotations of the ventricle around the apico-basal axis (the direction from the base of the heart to its apex), while permitting torsion. At the base $\Gamma^{\text{base}}_{0}$, an energy-consistent boundary condition is enforced: this ensures the correct stress distribution at the base boundary. For the endocardium $\Gamma^{\text{endo}}_{0}$, the applied boundary condition accounts for the pressure $p_{\text{LV}}(t)$, which reflects the pressure exerted by the blood within the ventricle. Lastly, the mechanical model $(\mathscr{M})$ influences $(\mathscr{A})$ by determining the local sarcomere length (SL); the sarcomeres are aligned with the muscle fibers, so the local sarcomere length directly correlates with the tissue stretch along the fiber direction.